Solve for x,y and z
Given:
x+10y=7
x+2z=4

Answer by april4869(19)

Solving for these variables is simply eliminating and dividing. For example...
The first equation, you need to solve for x and y, since z is not in the first equation.
x+10y=7
x=(7-10y)
And y:
10y=7-x
y=(7-x)/10
Now in this equation we need to solve for x and z, since y is not in this equation.
x+2z=4
x=(4-2z)
And z:
2z=4-x
z=(4-x)/2
Hope this helps:)